# Multi-Modal Optimization with k-Cluster Big Bang-Big Crunch Algorithm

### v1

@@ 8            $\textnormal{{P}}\leftarrow\textnormal{{evaluation}}(f,\textnormal{{P}})$;
II-E Time Complexity
We report the computational time complexity analysis of k-BBBC. As the population is divided into $k$ clusters, each cluster must perform a number of operations at least linear with the number of its individuals. However, since k-means is a stochastic algorithm, we cannot estimate the correct number of individuals per cluster that k-BBBC will feature throughout its run. Therefore, to evaluate the complexity of k-BBBC we will assume that each cluster contains a fixed number of individuals equal to $\nicefrac{{n}}{{k}}$ – which, as specified in (6), is $20$. This can be considered a good approximation over a statistically sufficient number of iterations of the algorithm.
@@ II-E Time Complexity
With the relation reported in (6), we can replace any occurrence of $n$ with the term $k$ (big-o notation ignores constants; thus, we can remove the coefficient $20$). The overall complexity of a single run of k-BBBC is then expressed as in (7), in which k-means features the dominant term. Note that this analysis is valid for both the elitist and non-elitist algorithms, as the number of individuals in the population is increased only by a constant factor of $k$ and the elites do not undergo a second re-evaluation.
@@
The complexity can also be expressed in terms of the problem’s properties, which include the parameters the number of optima $m$ and dimensionality $d$. By applying the formulation in (5), we obtain a complexity as defined in (8).
@@
III Post Processing
III-A Precise Identification of Detected Local Optima
k-BBBC produces as output a population of individuals that will very likely contain more solutions than the number of local optima $m$. It will return exactly $k$ solutions (i.e., the centers of mass), which must overestimate the actual number of local optima for complete retrieval as specified in Sec. II-C. This will cause local optima as well as other points in their proximity to appear in the output. In other words, although the population of k-BBBC converges to the local optima, it is still necessary to identify them within the output.
As shown in Fig. 1(c), k-BBBC converges to a well-separated population grouped around each local optima. Therefore, the simplest strategy to precisely identify the optimal solutions is using, once more, a clustering algorithm. The method described in Algorithm 3 proposes the following steps:
@@ III-A Precise Identification of Detected Local Optima
For this procedure, we preferred k-medoids to k-means due to its lower error rate as specified in Sec. II-C. Fig. 4 depicts the results of the identification method on the centers of mass retrieved by k-BBBC.
@@
III-B Quantifying the Number of Missed Optima
The procedure for optima identification described in Sec. III-A works well when the optimization method finds every local optimum in the search space. However, this is not always guaranteed: if the cardinality of the population $n$ is not large enough to explore the search space or if the number of clusters $k$ is insufficient, some local optima might be missed. When this happens, the $m$ identified solutions will contain repeated values as k-means is forced to divide the dataset into exactly $m$ clusters – similarly to the issue described in Fig. 2, one or more clusters will be split among different solutions. This is purely dependent on the user-defined parameters chosen while optimizing a specific problem, as well as on the number of generations employed during the execution – or sometimes, on the complexity of the problem that makes it difficult for some optima to be retrieved.
Suppose we have a problem with $m=4$ optima, and the optimization algorithm missed one. Then, the identification method will return four solutions, two of which being very close to each other. When observing this dataset, one would divide it into $k=3$ clusters rather than $4$ due to the proximity of the two solutions identifying the same optimum. Of course, also $k=4=m$ would work, as each data in the dataset can be considered a cluster itself; however, the best number of clusters would be $3$, indicating that the algorithm did miss one of them – Fig. 5 provides a graphical representation of this case. If the number of expected local optima $m$ is known, we then have a strategy to evaluate how many optima have been missed: observe the solutions retrieved by the identification method and determine the best number of clusters $k^{*}$ we could divide them into. If $k^{*}<m$, the optimization algorithm missed $m-k^{*}$ optima. Note that this process does not require knowing the optima’s location in the search space and can be used for any problem.
@@ III-B Quantifying the Number of Missed Optima
Figure 6: Examples of failure detection in the optimization methods. The plots on the left-hand side show the minimization of the Key function described in (1), with $m=24$; whereas the plots on the right-hand side show the silhouette analysis. We ran k-BBBC with $k=m$ in both examples to get an analyzable failure (both cases missed $8$ local optima). In the first example (a), k-BBBC was executed with the right parameters (as reported in Sec. II-C), and the silhouette features a clear maximum value for $k=24$, suggesting that no optima have been missed. In the second example (b), k-BBBC was executed with a higher number of generations (i.e., $g=350$); however, the value of $k=m$ was still insufficient to locate all local optima. Therefore, some of the $k$ solutions retrieved by k-BBBC are very close to each other, similar to what happens in Fig. 2 (not visible in the figure due to zoom). These very-close-to-each-other solutions can be grouped in a single cluster or can define their own cluster; hence, the plateau in the silhouette. The plateau ends at $k=16$, meaning that k-BBBC retrieved only $24-16=8$ local optima. In the third example (c), k-BBBC was executed with a small number of generations (i.e., $g=100$), which was not sufficient to crunch the individuals to the local optima; the silhouette features a plateau in the range $k=[18,24]$, and a peak on $k=16$ ($24-16=8$ missed optima): this is because the identification method (Sec. III-A) returned three points around the local optima at $0.66$, which can be divided into one cluster as well as into three clusters.
@@
    When the optimization algorithm correctly retrieves $m_{c}=m$ local optima, the silhouette values feature a maximum point at $k^{*}=m$. This is shown in Fig. 6(a).
@@
    When the optimization algorithm misses a total of $m_{m}$ local optima, the silhouette score for all the values of $k\in[m-m_{m},m]$ will be close to $1$, as shown in Fig. 5. This results in a plateau of silhouette scores that represent different correct clusterizations of the given dataset; however, the value of $k^{*}$ is found at the left corner of the plateau. In Fig. 6(b), a total of $m_{m}=8$ over $m=24$ optima were missed, and the silhouette trend features a plateau ending at $k^{*}=16$, resulting in $m-k^{*}=24-16=8=m_{m}$ missed optima and $k^{*}=16=m_{c}$ correct identified optima.
@@
    When the optimization algorithm misses a total of $m_{m}$ local optima due to poor convergence, some values can be randomly scattered through the search space as shown in Fig. 6(c). In this case, the silhouette scores will feature the plateau preceded by a peak point, which is the value of $k^{*}$. However, this case is considered unstable and prone to detection mistakes: waiting for the optimization algorithm to converge correctly is recommended to avoid such occurrences. Note that the peak is visible only because the three solutions at $x=0.6455$ are sufficiently close to each other to be considered as a cluster.
@@ 6                   $\textnormal{{m}}_{c}\leftarrow k$;
9      $\textnormal{{m}}_{m}\leftarrow\textnormal{{m}}-\textnormal{{m}}_{c};$ return $\textnormal{{m}}_{c},\textnormal{{m}}_{m}$;
10
@@
IV Experiments
In this section, we report the results of k-BBBC (both elitist and not) compared to another method in the literature: Multimodal Cuckoo Search (MCS) [6]. We selected only this algorithm for two reasons: (i) based on the original work, it outperforms other multi-modal algorithms such as Crowding Differential Evolution [37], Fitness Sharing Differential Evolution [30, 37], Clearing Procedure [31], Elitist-Population Strategy [26], Clonal Selection Algorithm [55], and Artificial Immune Network [56] – so we did not need to implement each of them for testing; and (ii) its code is publicly available on MathWorks File Exchange.
For k-BBBC, we ran the post-processing described in Sec. III – that is because k-BBBC ensures that the entire population converges to the optima. However, for MCS we ran a different post-processing, since it does not ensure the same trend: after termination, we executed its depuration procedure [6] to identify the local optima within the retrieved population – similarly to our identification procedure described in Sec. III-A. This should return a number of individuals that is $<=m$; however, in case this number is larger, we executed our identification procedure. Lastly, this dataset of individuals goes through the quantification method of Sec. III-B to assess how many optima have been missed.
IV-A Methodology
The algorithms were compared on the twenty low-dimensionality (i.e., $d=1$ or $2$) and four high-dimensionality problems (i.e., $d>4$). We executed every algorithm with the same population size, defined as in (5) and (6) and based on the problem to be optimized. To ensure a fair comparison, we executed k-BBBC for $1000$ generations, recorded its runtime for each function, and set it as a stopping condition for MCS (i.e., we stopped MCS at the end of the iteration that exceeded k-BBBC runtime – specifically, the average runtime of non-elitist k-BBBC). Although the authors did not report it, we observed that MCS has a higher time complexity than k-BBBC due to its depuration procedure; therefore, setting a fixed number of population and generations would have resulted in an unfeasible runtime for problems with a large number of optima. Every problem was solved $25$ times per algorithm, and every result is presented as the average and standard deviation of these runs.
@@ IV-B Evaluation Metrics
IV-C Experiments with low dimensionality problems
The twenty problems with low dimensionality are summarized in Table IV. For each problem, the exact local optima z were manually retrieved by running fmincon (Interior Point Algorithm [57]) on the search space’s sub-regions surrounding them – i.e., we manually located each concave region by observing the function’s graph and set its geometrical limits as bounds of the problem.
TABLE II: Results for Low Dimensionality Problems
[…]
IV-D Experiments with high dimensionality problems
We tested k-BBBC on the high-dimensional multimodal test problem defined in (11). This function is similar to Rastrigin’s function and it is a modified version of the Key function defined in (1). In the range $\left[0,1\right]$ and with the $j$ values defined as in (11), it features $48$ local minima (one of them being global) – the exact optima values z are reported in the work of Deb and Saha [7]. We tested the algorithms for different dimensionalities: $d=4$, $8$, $16$, and $32$. Due to the multi-dimensionality of the function, we will not report its graph.
subject to $0\leq x_{i}\leq 1,\quad i=1,2,...,d$
TABLE III: Results for High Dimensionality Problems
[…]
MCS -- -- -- -- -- -- --
(a) Accuracy (Search Space)
(b) Success Rate
IV-E Results
Table II and III show the results for problems having low and high dimensionality, respectively – reporting average and standard deviation for each metric.
The plot in Fig. 7(a) reports the overall accuracy in search space for each algorithm, and each result was normalized against their function’s domain. The plot evidences a higher accuracy of elitist k-BBBC, outperforming both k-BBBC and MCS. Note that the accuracy in objective space was not reported as a plot due to the challenge of normalizing values against the codomain of each problem. Additionally, Table III shows that k-BBBC and its elitist version feature a higher accuracy (in both search and objective space) compared to MCS; more importantly, MCS could not find any optima for problems with $d\geq 8$, and could not finish its execution with $d=32$ due to extensive computational requests, suggesting that k-BBBC can tackle high dimensionality problems better than other algorithms.
Fig. 7(b) summarizes the actual success rate of each algorithm (i.e., retrieved solutions compared against the actual local optima). Although MCS performs fairly well in most of the problems, it is outperformed by both versions of k-BBBC – this is evident in $f_{6}$, $f_{8}$, $f_{9}$, $f_{15}$, and $f_{16}$. Furthermore, as previously mentioned, MCS could not find any optima for functions with $d\geq 8$, resulting in a success rate equal to 00.
As the runtime was set to be consistent for all algorithms, we also reported the number of function evaluations during each run. Both k-BBBC and its elitist version feature the same number of evaluations, which is constant throughout all the runs as individuals are only evaluated once per generation (note that the elites are not re-evaluated in following generations); on the other hand, MCS features a variable number of evaluations due to its depuration procedure. All algorithms show a similar number of evaluations, with MCS performing generally fewer evaluations depending on when the depuration procedure is triggered; this, however, might affect the results of the optimization – i.e., if the run is terminated abruptly, MCS might not have had the time to retrieve neighbors and more optimal solutions. This might also suggest an alternative observation, which is MCS has a higher time complexity than k-BBBC (although the authors did not report any analysis in the original work [6]), leading to a termination that could also happen midrun and therefore preventing the algorithm to converge to all local optima.
V Discussion
Based on the results reported in Sec. IV, we can summarize several advantages and disadvantages of k-BBBC.
V-A Advantages
We report a list of the advantages of k-BBBC:
  • The algorithm features high accuracy in both search space and objective space while retrieving (almost) all the local optima for every problem – regardless of their dimensionality. This is especially evident while using elitism.
  • In most cases, k-BBBC outperforms MCS in terms of accuracy and success rate, and therefore the other multi-modal optimization algorithms that were outperformed by MCS (as reported in the original work by Cuevas and Reyna-Orta [6]).
  • The entire population converges to the local optima, not only the best individuals. This is due to the denominator term in (4), which reduces the range of exploration of new individuals from their center of mass. The range of exploration is decreased at each generation of k-BBBC (and equivalently to its original version for global optimization), guaranteeing convergence in the entire population. Having the entire population converging to the local optima allows for easy identification of the best individuals, and inspired us to develop the procedures described in Sec. III.
  • Directly from the previous advantage, using the quantification method (Algorithm 4) allows us to have an estimation of how the algorithm performed in terms of success rate without the need of knowing the exact value of the actual optima – which we will not have because the purpose of a multi-modal optimization algorithm is to find those values. Based on the results summarized in Sec. IV-E, where we compared the actual success rate with the one obtained by quantification, we can claim that the quantification method works properly when the population converges to local optima. This, however, can be exploited only if the number of optima $m$ is known or can be estimated.
  • The algorithm can tackle high-dimensional problems with full retrieval of all local optima (see Table III), whereas MCS failed. Additionally, when compared with the work by Deb and Saha [7], their bi-objective NSGA-II-based algorithm was not tested for problems with $d>16$, whereas k-BBBC works well also for $d=32$.
  • k-BBBC works for both minimization and maximization problems.
V-B Disadvantages
We report a list of the disadvantages of k-BBBC:
  • It is required to know the number of local optima of a problem or to have a valid estimation of it. This might not always be possible; however, this number can be empirically retrieved: one could get a population that correctly converges to all local optima; then, use the silhouette or even manually count the number of optima, and use Algorithm 3 to find out the actual number.
  • It relies on a mathematical estimation of the number of clusters for correct retrieval. Although the formulation reported in (5) and (6) are appropriate for the problems tested in this study, it is not a guarantee of full reliability.
  • It is based on clustering, and therefore, relies on algorithms such as k-means and k-medoids, with all their related issues and limitations (e.g., see Fig. 2). Furthermore, when running k-BBBC on the problem with $d=32$, we observed that k-means failed to converge in 200 iterations (a limit we empirically set based on the problems we tested). This did not prevent k-BBBC from retrieving every optimum, but it is important to report this issue – to the best of our knowledge, there is no hard limit for the dimensionality that k-means can tackle; however, it is reasonable to assume that the higher the dimensionality, the less the convergence.
  • Problems featuring non-uniform distribution of local optima in the search space (e.g., Schwefel $f_{8}$ and $f_{9}$, Himmelblau $f_{10}$, and Vincent $f_{15}$) may deceive the clustering algorithm, especially in regions with a high density of local optima – i.e., k-means can group all those solutions in the same cluster. However, the problem is found also for MCS, with k-BBBC outperforming it in each of the aforementioned problems.
  • Based on the analysis reported in Sec. II-E, the algorithm’s runtime increases quadratically with the number of local optima of the problem and cubically with its dimensionality. Therefore, problems featuring high dimensionality are expected to have a higher runtime – e.g., as shown in Table III, a single run for $d=32$ took approximately eleven hours. However, we can consider optimization – in most cases – as an offline process, and this timing still falls within an acceptable range.
  • The algorithm cannot distinguish a local optimum from a point lying on a plateau: this is what happened to Bird function $f_{11}$, in which sometimes a plateau point was mistaken for a local optimum. However, this problem is visible also in MCS, and it can be easily fixed by increasing the number of local optima expected $m$. In our trials, we observed that a value of $m=6$ (two more than required) was sufficient to correctly retrieve all four optima and two random plateau points.
VI Conclusion
In this paper, we presented k-BBBC, an MMEA based on clustering. Additionally, we introduce a post-processing procedure to identify local optima in a population of solutions and quantify how many have been missed by the algorithm without the need to compare them with the actual points. We compared k-BBBC with other algorithms in the literature (both directly and indirectly), and our results highlight high accuracy and, particularly, a higher success rate in retrieving optima. Additionally, k-BBBC can tackle problems with a high number of optima (tested on $m=379$) and high dimensionalities (tested on $d=32$).
In future works, we aim to address some of the disadvantages of k-BBBC, such as: (i) detecting when a specific cluster is converging to a plateau rather than a local optimum, and (ii) proposing alternative algorithms to remove the burden of knowing the number of expected optima. Lastly, we plan to test the algorithm on real-engineering applications, such as retrieving alternative sub-optimal designs for robotic devices [1].
Code
The MATLAB script is available on EVO Lab’s MathWorks File Exchange repository.
Appendix: Benchmark Low-Dimensional Problems
Table IV reports the list of twenty benchmark problems used in the experiment of Sec. IV-C.
TABLE IV: Benchmark Problems (1/3)

### v2

@@ 9            $\textnormal{{P}}\leftarrow\textnormal{{evaluation}}(f,\textnormal{{P}})$;
2.5 Time Complexity
We report the computational time complexity analysis of k-BBBC. As the population is divided into $k$ clusters, each cluster must perform a number of operations at least linear with the number of its individuals. However, since k-means is a stochastic algorithm, we cannot estimate the correct number of individuals per cluster that k-BBBC will feature throughout its run. Therefore, to evaluate the complexity of k-BBBC we will assume that each cluster contains a fixed number of individuals equal to $\nicefrac{{n}}{{k}}$ – which, as specified in Eqn. (5), is $20$. This can be considered a good approximation over a statistically sufficient number of iterations of the algorithm.
@@ 2.5 Time Complexity
With the relation reported in Eqn. (5), we can replace any occurrence of $n$ with the term $k$ (big-o notation ignores constants; thus, we can remove the coefficient $20$). The overall complexity of a single run of k-BBBC is then expressed as in Eqn. (6), in which k-means features the dominant term. Note that this analysis is valid for both the elitist and non-elitist algorithms, as the number of individuals in the population is increased only by a constant factor of $k$ and the elites do not undergo a second re-evaluation.
@@
The complexity can also be expressed in terms of the problem’s properties, which include the parameters number of optima $m$ and dimensionality $d$. By applying the formulation in Eqn. (4), we obtain a complexity as defined in Eqn. (7).
@@
3 Post Processing Methods
3.1 Precise Identification of Detected Local Optima
k-BBBC produces as output a population of individuals that will very likely contain more solutions than the number of local optima $m$. It will return exactly $k$ solutions (i.e., the centers of mass), which must overestimate the actual number of local optima for complete retrieval as specified in Sec. 2.3. This will cause local optima as well as other points in their proximity to appear in the output. In other words, although the population of k-BBBC converges to the local optima, it is still necessary to identify them within the output.
As shown in Figure 1(c), k-BBBC converges to a well-separated population grouped around each local optima. Therefore, the simplest strategy to precisely identify the optimal solutions is using, once more, a clustering algorithm. The method described in Algorithm 3 proposes the following steps:
  • 1.
@@ 3.1 Precise Identification of Detected Local Optima
For this procedure, we preferred k-medoids to k-means due to its lower error rate as specified in Sec. 2.3. Figure 4 depicts the results of the identification method on the centers of mass retrieved by k-BBBC.
@@
3.2 Quantifying the Number of Missed Optima
The procedure for optima identification described in Sec. 3.1 works well when the optimization method finds every local optimum in the search space. However, this is not always guaranteed: if the cardinality of the population $n$ is not large enough to explore the search space or if the number of clusters $k$ is insufficient, some local optima might be missed. When this happens, the $m$ identified solutions will contain repeated values as k-means is forced to divide the dataset into exactly $m$ clusters – similarly to the issue described in Figure 2, one or more clusters will be split among different solutions. This is purely dependent on the user-defined parameters chosen while optimizing a specific problem, as well as on the number of generations employed during the execution – or sometimes, on the complexity of the problem that makes it difficult for some optima to be retrieved.
@@ 3.2 Quantifying the Number of Missed Optima
Figure 6: The plots show the minimization of the Key function with $m=24$ along with its silhouette analysis. In (a), k-BBBC was executed with $k=2m$ and the silhouette features a clear maximum value for $k=24$ – no optima have been missed. In (b), k-BBBC was executed with $k=m$, which was insufficient to locate all local optima. Therefore, some of the $k$ solutions retrieved by k-BBBC are very close to each other (like in Figure 2). These solutions can be grouped in a single cluster; hence, the plateau in the silhouette ending at $k=16$ – k-BBBC retrieved only $24-16=8$ local optima. In (c), k-BBBC was executed with $k=m$ and ran for fewer generations than in (b), which was not sufficient to crunch the individuals to the local optima. Its silhouette features a plateau in the range $k=[18,24]$, and a peak on $k=16$ ($24-16=8$ missed optima): this is because the identification method (Sec. 3.1) returned three points around the local optima at $0.66$, which can be divided into one cluster as well as into three clusters.
Suppose we have a problem with $m=4$ optima, and the optimization algorithm missed one. Then, the identification method will return four solutions, two of which being very close to each other. When observing this dataset, one would divide it into $k=3$ clusters rather than $4$ due to the proximity of the two solutions identifying the same optimum. Of course, also $k=4=m$ would work, as each data in the dataset can be considered a cluster itself; however, the best number of clusters would be $3$, indicating that the algorithm did miss one of them – Figure 5 provides a graphical representation of this case. If the number of expected local optima $m$ is known, we then have a strategy to evaluate how many optima have been missed: observe the solutions retrieved by the identification method and determine the best number of clusters $k^{*}$ we could divide them into. If $k^{*}<m$, the optimization algorithm missed $m-k^{*}$ optima. Note that this process does not require knowing the optima’s location in the search space and can be used for any problem.
@@
    When the optimization algorithm correctly retrieves $m_{c}=m$ local optima, the silhouette values feature a maximum point at $k^{*}=m$. This is shown in Figure 6(a).
@@
    When the optimization algorithm misses a total of $m_{m}$ local optima, the silhouette score for all the values of $k\in[m-m_{m},m]$ will be close to $1$, as shown in Figure 5. This results in a plateau of silhouette scores that represent different correct clusterizations of the given dataset; however, the value of $k^{*}$ is found at the left corner of the plateau. In Figure 6(b), a total of $m_{m}=8$ over $m=24$ optima were missed, and the silhouette trend features a plateau ending at $k^{*}=16$, resulting in $m-k^{*}=24-16=8=m_{m}$ missed optima and $k^{*}=16=m_{c}$ correct identified optima.
@@
    When the optimization algorithm misses a total of $m_{m}$ local optima due to poor convergence, some values can be randomly scattered through the search space as shown in Figure 6(c). In this case, the silhouette scores will feature the plateau preceded by a peak point, which is the value of $k^{*}$. However, this case is considered unstable and prone to detection mistakes: waiting for the optimization algorithm to converge correctly is recommended to avoid such occurrences. Note that the peak is visible only because the three solutions at $x=0.6455$ are sufficiently close to each other to be considered as a cluster.
@@ 7                   $\textnormal{{m}}_{c}\leftarrow k$;
9
10      $\textnormal{{m}}_{m}\leftarrow\textnormal{{m}}-\textnormal{{m}}_{c};$
11      return $\textnormal{{m}}_{c},\textnormal{{m}}_{m}$;
12
@@ 10      $\textnormal{{m}}_{m}\leftarrow\textnormal{{m}}-\textnormal{{m}}_{c};$
4 Experiments for Finding Local Peaks (on Basic Multimodal Problems)
In this section, we report the results of k-BBBC (both elitist and not) compared to another method in the literature: Multimodal Cuckoo Search (MCS) [7]. We selected only this algorithm for two reasons: (i) based on the original work, it outperforms other multimodal algorithms such as CDE [38], FSDE [31, 38], CP [32], EPS [27], CSA [57], and AIN [58] – so we did not need to implement each of them for testing; and (ii) its code is publicly available on MathWorks File Exchange.
For k-BBBC, we ran the post-processing described in Sec. 3 – that is because k-BBBC ensures that the entire population converges to the optima. However, for MCS we ran a different post-processing, since it does not ensure the same trend: after termination, we executed its depuration procedure [7] to identify the local optima within the retrieved population – similarly to our identification procedure described in Sec. 3.1. This should return a number of individuals that is $<=m$; however, in case this number is larger, we executed our identification procedure. Lastly, this dataset of individuals goes through the quantification method of Sec. 3.2 to assess how many optima have been missed.
4.1 Methodology
The algorithms were compared on the twenty low-dimensionality (i.e., $d=1$ or $2$) and four high-dimensionality problems (i.e., $d>4$). We executed every algorithm with the same population size, defined as in Eqn. (4) and (5) and based on the problem to be optimized. To ensure a fair comparison, we executed k-BBBC for $1000$ generations, recorded its runtime for each function, and set it as a stopping condition for MCS (i.e., we stopped MCS at the end of the iteration that exceeded k-BBBC runtime – specifically, the average runtime of non-elitist k-BBBC). Although the authors did not report it, we observed that MCS has a higher time complexity than k-BBBC due to its depuration procedure; therefore, setting a fixed number of population and generations would have resulted in an unfeasible runtime for problems with a large number of optima. Every problem was solved $25$ times per algorithm, and every result is presented as the average and standard deviation of these runs.
@@ 4.2 Evaluation Metrics
4.3 Experiments with low dimensionality problems
Table 2: Basic Low-Dimension Multimodal Problems